# Verbosity-Aware Rationale Reduction: Effective Reduction of Redundant Rationale via Principled Criteria

### v2

@@ 4.3 Contrasting with Wrong Answer
Inspired by the miscalibrated log-likelihoods between accepted and rejected responses during standalone Supervised Fine-Tuning (SFT) in alignment learning Rafailov et al. (2024); Azar et al. (2024); Hong et al. (2024), we also assess whether the reduced rationales $R^{\prime}$ result in inaccuracy answer generation by introducing a wrong answer $y_{w}$. Instead of employing the correct answer distribution $q(y|x)$ in Equation (2), we initiate our formula with the wrong answer distribution $q^{\prime}$ (i.e., $1-q(y|x)$, normalized sum to 1) Through algebraic manipulation, we can obtain:
@@
where $\{y_{w}^{(k)}\}_{k\in\left[K\right]}$ is sampled from the in-batch negatives depending on the dataset. Consequently, the $\textit{verbosity}(y_{w})$ is computed as:
@@
Since computational constraints necessitate sampling incorrect answers to calculate $\textit{verbosity}(y_{w})$, we contrast $\textit{verbosity}(y_{w})$ with $\textit{verbosity}(y_{g})$ rather than solely using $\textit{verbosity}(y_{w})$. This ensures that the likelihood of incorrect answers remains lower than that of ground truth answers, even after the removal of $r_{i}$ as follows:
@@
A smaller $\textit{verbosity}(y_{w})$ value compared to $\textit{verbosity}(y_{g})$ indicates that $R^{\prime}$ increases the likelihood of generating incorrect answers over the ground truth answer after removing $r_{i}$, making it beneficial to remove.
@@ 4.4 CoT Training with Rationale Reduction
In our framework, the model is trained using Equation (4) for predefined warm-up steps to inject its reasoning capabilities. After the warm-up steps, at each training step $t$, each sentence $r_{i}$, starting from the first index sentence in $R$, is evaluated against either Equation (5) alone (denoted as VARR) or both Equations (5) and (9) (denoted as VARR+). Sentences satisfying these respective criteria are selected for removal and are not included in subsequent training steps. The maximum removable number of sentences at each training step $t$ is determined based on a linear schedule, adopting Deng et al. (2024b)’s setting as follows:
@@
where $T$ represents the total number of training steps, $N_{t}$ is the total number of rationale sentences at step $t$, and $r(t)$ indicates the maximum number of sentences that can be removed at that step.
@@ 5.1 Training Configuration
Figure 3: Performance for 5 different reasoning tasks. Acc refers to Accuracy while Tokens denotes the Generated Tokens. Bar plots illustrate the accuracy of each method on datasets, while line plots show the average generated tokens by each method across the datasets. The y-axis represents the range of each metric, and the x-axis lists the datasets evaluated.
@@
We conducted experiments across two categories to provide a comprehensive evaluation of VARR’s versatility and effectiveness, in contrast to prior research that predominantly focuses on simple arithmetic tasks like multi-digit multiplication (Deng et al., 2023, 2024b). Initially, we evaluate with arithmetic reasoning tasks, including datasets like GSM8K (G8K; Cobbe et al. 2021a) and MathQA (MQA; Amini et al. 2019a). We also examine the performance of our method on commonsense reasoning tasks, employing datasets including CommonsenseQA (CQA; Talmor et al. 2019), TrivaiQA (TQA; Joshi et al. 2017), and StrategyQA (SQA; Geva et al. 2021).
@@
The warm-up stage is set to 0.1 of the total training steps, and its impact is analyzed in Section 5.5.4. In addition, an optimizer is reinitialized at the beginning of each epoch to stabilize the model training inspired by Deng et al. (2024b)’s setting, with its effects described in Section 5.5.3. All baselines and our proposed method are trained for 5 epochs to ensure fairness in comparison. Detailed implementation of VARR/VARR+ can be found in the Appendix B.
@@ 5.2 Baselines
We compare our method to the following baselines: Explicit-CoT (Nye et al., 2021), where the model is finetuned with explicit chain-of-thought reasoning, and ICoT-SI (Deng et al., 2024b), where the model is fine-tuned using a linear token elimination schedule. It is important to note that after removing intermediate reasoning paths, the model’s performance should be competitive with Explicit-CoT. Consequently, ICoT-SI serves as our primary baseline. We exclude Implicit-CoT Deng et al. (2023) from our baselines due to its significant memory requirements, requiring the training of three models on a single GPU. Furthermore, according to Deng et al. (2024b), this method underperformed compared to ICoT-SI. All baselines were trained on a single A100-80GB GPU, and detailed training configurations for each method are provided in Appendix C.
@@ 5.4 Main results
In Figure 3, we present the results for each reasoning task across different models. Our method demonstrates consistent performance improvements or maintains comparable results across most benchmarks, while also achieving more efficient reasoning path generation. Notably, we observed significant improvements in average accuracy with increments of 2.59%, 6.27%, and 14.27% for Mistral 7B, Llama3.2 1B, and Llama3.2 3B (VARR+), respectively, compared to Explicit CoT, while efficiency improved by 19.35%, 17.24%, and 14.88% for these models. These findings indicate that some reasoning sentences are redundant and can even disturb answer generation, highlighting that effective reasoning capabilities can be learned through selective utilization of essential reasoning units.
For ICoT-SI, the model exhibited performance degradation after token removal training. Specifically, ICoT-SI shows an average performance degradation of 18.93% compared to Explicit CoT. Furthermore, while the Mistral 7B model achieved a 50.26% accuracy on the GSM8K using a large augmented dataset with succinctly generated reasoning paths (trained with the 385K-size dataset used in the original paper (Deng et al., 2024b)), its performance dropped significantly to 28.27% when trained on the standard GSM8K dataset (trained with the 7K-size dataset). This indicates that while ICoT-SI may excessively reduce the length of rationales in reasoning tasks, identifying the appropriate reasoning unit and applying principled criteria are crucial for maintaining performance in real-world application problems. Qualitative results across baselines are presented in Appendix F, and a detailed analysis of the reasoning unit will be discussed in Section 5.5.1.
Moreover, introducing a wrong answer in VARR led to performance improvements in most datasets. While solely using the verbosity($y_{g}$) enhanced efficient generation, it is notable that verbosity($y_{w}$) constraints reduced rationales to increase the likelihood of generating incorrect answers, which contributes to the robustness and stability of the training process.
@@ 5.5.1 Identifying Appropriate Units for Removal
Table 1: Analysis across various reasoning reduction units and the application of principled criteria. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
In this section, we experimentally demonstrate why the sentences are suitable units for reduction reasoning units. Unlike ICoT-SI which lacks specific criteria for token removal, we apply VARR+ at the token level (denoted as VARR-Tok) to assess whether tokens can effectively function as reduction units in reasoning processes. As depicted in Table 1, applying principled criteria at the token level results in an average performance gain of 26.95%, indicating that the principled criteria contribute to robust performance. Moreover, extending the reasoning units from tokens to sentences (denoted as VARR-Sent) leads to a further performance increase of 21.83% over VARR-Tok, highlighting the importance of selecting the appropriate reasoning unit for effective reasoning path reduction. These findings demonstrate the importance of selecting appropriate reasoning units, with sentences serving as linguistically natural and effective boundaries for the removal process.
@@ 5.5.2 Analyzing the Impact of Sentence Position on Removal Efficacy
Table 2: Performance across 5 different reasoning tasks, evaluated with different sentence position removal. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
@@
In Section 3, we demonstrated that gradually removing sentences from random and back positions can degrade model performance. To further explore this finding and assess the robustness of removing sentences from the front position, we conducted experiments with unguided random sentence removal (denoted as No Rule) and applied VARR+ with random position and reverse sentence order (denoted as Random and Back, respectively). As shown in Table 2, unguided random sentence removal resulted in a 25.30% decrease in performance relative to our method, highlighting the critical role of systematic configuration even after selecting sentences as the units of reduction. Additionally, Random and Back strategies result in an average 7.50% performance degradation compared to our method. This finding again indicates that earlier sentences in the reasoning path tend to be more redundant, and prioritizing their removal effectively balances rationale reduction while maintaining reasoning performance.
5.5.3 Reinitializing the Optimizer
G8K MQA CQA TQA SQA
w/o reinit 45.18 48.70 87.82 83.72 76.74
(115.91) (92.86) (32.95) (38.27) (33.89)
w/ reinit 54.98 56.95 89.56 83.45 78.19
(100.38) (91.04) (33.55) (39.17) (35.12)
Table 3: Performance Comparison before and after the reinitializing optimizer in every training epoch. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
We reinitialized the optimizer after each training epoch to stabilize training, inspired by Deng et al. (2024b). We employed the AdamW Loshchilov and Hutter (2019) optimizer, where the first and second moments are gradually updated based on current gradients, which could lead to unstable training in our framework. Specifically, after applying VARR throughout the dataset, a reduction in the rationale for some data between epochs could lead to instability in the training process. With a 19.32% performance gain achieved by reinitializing the optimizer, we reconfirm that Deng et al. (2024b)’s findings are robust across diverse datasets with complex semantic and syntactic structures in reasoning paths.
5.5.4 Varying the Warm-up Steps
@@
Figure 4: Accuracy and Generated Token difference across various warm-up stages.
@@
We evaluated our method against various warm-up stages on the TriviaQA and CommonsenseQA datasets. As the duration of the warm-up stages increases, the model becomes more fitted to the non-reduction dataset, which inhibits VARR’s ability to eliminate redundant sentences from the reasoning path. Consequently, as illustrated in Figure 4, longer warm-up periods result in an increase in generated tokens and a decrease in accuracy. These results suggest that 0.1 training steps provide sufficient time to inject reasoning abilities while enabling the systematic reduction of redundant reasoning sentences during the learning process. Therefore, we set 0.1 as the default setting for the warm-up steps.
@@ 5.5.5 Removal Ratio Analysis
Figure 5: Removal ratio of redundant sentences throughout the training stage. The Removal Ratio, indicated on the y-axis, is calculated by dividing the number of sentences actually removed by the maximum potential removal ratio (i.e., ${size(\{r_{i}\})/r(t)})$. The x-axis represents the index of the epochs.
In Figure 5, we analyze the actual amount of rationale sentences removed during training. We examine it by calculating the removal ratio, the proportion of actual removed sentences to the maximum potential removal sentences $r(t)$. Our analysis indicates that not all sentences designated for maximum removal range are always eliminated during the training process. Notably, a significant proportion of redundant sentences are removed in the early stages of training, with fewer sentences being removed as the model progresses through the middle to later stages, thereby stabilizing its training. A similar trend is observed across other datasets, as detailed in Appendix E.
@@ 6 Conclusion
In this work, we propose sentence-level rationale reduction framework VARR and provide empirical evidence that models trained with non-redundant rationales achieve greater efficiency. We address the lack of principled criteria for identifying redundant sentences during the training phase by developing a reduction method that not only preserves the model’s reasoning capabilities but also reduces the likelihood of generating incorrect answers. Our experiments show that VARR can efficiently handle a diverse range of tasks with fewer generated tokens, without sacrificing its accuracy. Further empirical findings confirm that sentences are appropriate reasoning units for rationale reduction, highlighting the necessity of principled criteria for robust application. This work contributes novel insights to rationale reduction research, contributing to the efficient elicitation of reasoning in language models.
@@
While our work provides novel insights into rationale reduction research, our experiments were primarily conducted using a relatively small large language model, constrained by computational costs. Additionally, for the same reasons, it was not feasible to test the model with datasets featuring long sequences in both queries and rationales (Reddy et al., 2024; Yu et al., 2024a). However, given that the VARR/VARR+ frameworks are designed with a systematically structured approach, we believe that they could also be effective in scaled-up settings.

### v3

@@ 4.3 Contrasting with Wrong Answer
Inspired by the miscalibrated log-likelihoods between accepted and rejected responses in standalone Supervised Fine-Tuning (SFT) for alignment learning Rafailov et al. (2024); Azar et al. (2024); Hong et al. (2024), we examine whether reduced rationales $R^{\prime}$ lead to inaccurate answer generation by incorporating a wrong answer $y_{w}$.
Instead of employing the correct answer distribution $q(y|x)$ in Equation (2), we initiate our formula with the wrong answer distribution $q^{\prime}$ (i.e., $1-q(y|x)$, normalized sum to 1). Through algebraic manipulation, we derive:
@@
where $\{y_{w}^{(k)}\}_{k\in\left[K\right]}$ is sampled from the in-batch negatives depending on the dataset. Consequently, $\textit{verbosity}(y_{w})$ is computed as:
@@
Since computational constraints necessitate sampling incorrect answers to calculate $\textit{verbosity}(y_{w})$, we evaluate the effectiveness of removal by comparing $\textit{verbosity}(y_{w})$ against $\textit{verbosity}(y_{g})$ rather than solely using $\textit{verbosity}(y_{w})$ as follows:
@@
When both conditions $\textit{verbosity}(y_{g})\geq 0$ and $\textit{verbosity}(y_{w})-\textit{verbosity}(y_{g})\leq 0$ are satisfied, it indicates that removing $r_{i}$ from $R$ not only improves the model’s performance but also increases its preference for the ground truth answer over incorrect answers, supporting the removal of $r_{i}$.
@@ 4.4 CoT Training with Rationale Reduction
In our framework, the model is trained using Equation (4) for predefined warm-up stage to inject its reasoning capabilities. Subsequently, at each training step $t$, we evaluate each sentence $r_{i}$ sequentially from the first sentence in $R$, using either Equation (5) alone (denoted as VARR) or the combination of Equations (5) and (9) (denoted as VARR+). Sentences that satisfy these respective criteria are identified for removal and excluded from subsequent training steps. The maximum removable number of sentences at each training step $t$ is determined based on a linear schedule, adopting ICoT-SI (Deng et al., 2024b)’s setting as follows:
@@
where $T$ represents the total number of training steps, $N_{t}$ is the total number of rationale sentences at step $t$, and $r(t)$ indicates the maximum number of sentences that can be removed at that step. Note that unlike ICoT-SI, which enforcely removes a predefined number of tokens during training, our method preserves the essential reasoning steps by employing principled removal criteria.
@@ 5.1 Training Configuration
Figure 3: Pareto plot of accuracy versus the number of generated tokens. The gray dotted lines connect the Pareto frontiers of the baselines, and our VARR (or VARR+) consistently outperforms the pareto frontiers across all subplots. While ICoT-SI and Coconut substantially trade-off accuracy for efficiency, VARR/VARR+ maintains high accuracy while reducing generated tokens, demonstrating its superior efficiency-performance balance.
@@
We conducted experiments across two categories to provide a comprehensive evaluation of VARR’s versatility and effectiveness, in contrast to prior research that predominantly focuses on simple arithmetic tasks like multi-digit multiplication (Deng et al., 2023, 2024b). Initially, we evaluate with arithmetic reasoning tasks, including datasets like MathQA (MQA; Amini et al. 2019a) and GSM8K (G8K; Cobbe et al. 2021a). We also examine the performance of our method on commonsense reasoning tasks, employing datasets including CommonsenseQA (CQA; Talmor et al. 2019), TriviaQA (TQA; Joshi et al. 2017), and StrategyQA (SQA; Geva et al. 2021). Note that unlike previous works (Deng et al., 2023, 2024b), we do not synthesize training data (especially the intermediate steps) to validate the generalizability and applicability of our method.
@@
The warm-up stage is set to 0.1 of the total training steps, and its impact is analyzed in Section 5.5.3. In addition, an optimizer is reinitialized at the beginning of each epoch to stabilize the model training inspired by Deng et al. (2024b)’s setting, with its effects described in Appendix E. All methods are trained for 5 epochs for fair comparison, with detailed implementation of VARR/VARR+ provided in Appendix B.
@@ 5.2 Baselines
We compared our method against the following baselines: Explicit-CoT (Nye et al., 2021), where the model is finetuned with explicit chain-of-thought reasoning; ICoT-SI (Deng et al., 2024b), where the model is fine-tuned using a linear token elimination schedule; and Coconut (Hao et al., 2024), where the model is fined-tuned to compress rationales into a predefined number of token hidden states. We excluded Implicit-CoT Deng et al. (2023) from our evaluation due to its substantial computational demands, specifically requiring three models to be trained simultaneously on a single GPU. Moreover, Deng et al. (2024b) demonstrated that this method achieves a lower performance compared to ICoT-SI. Given that both our baselines and VARR aim to maintain Explicit-CoT’s performance while reducing the number of generated tokens, we establish Explicit-CoT’s performance as our primary baseline for comparison. All baselines were trained on a single A100-80GB GPU, and detailed training configurations for each method are provided in Appendix C.
@@ 5.4 Main results
In Figure 3, we present the results for each reasoning task across different models. VARR/VARR+ achieves comparable or superior performance compared to Explicit-Cot across most datasets while reducing the average token generation. Specifically, VARR+ significantly increases performance by an average of 7.71% across all datasets and models, while improving efficiency by reducing token generation by 19.37% on average.
It is noteworthy that these findings contrast with ICoT-SI, demonstrating that performance can be improved while reducing the number of generated tokens. This suggests that existing reasoning data contains unnecessary reasoning sentences that may harm performance. Furthermore, effective reasoning can be achieved by selectively removing sentences based on appropriate criteria.
However, the baselines exhibit performance degradation compared to Explicit-CoT. For ICoT-SI, we observe an average performance decline of 21.98%, while Coconut shows a degradation of 25.20%, demonstrating their imbalanced trade-off in efficiency. These results suggest that heuristic reasoning reduction approaches do not effectively induce implicit reasoning within the model as discussed by ICoT-SI and Coconut. This indicates that identifying and retaining appropriate reasoning units through principled criteria is crucial for maintaining performance in practical applications. A more detailed discussion of the impact of the choice of reduction unit and criterion is provided in Section 5.5.1.
Furthermore, qualitative analysis (refer to Appendix J) confirms that ICoT-SI and Coconut fail to generate valid reasoning paths for answer generation, while VARR+ produces concise, yet effective reasoning paths that lead to correct answers. Additionally, the incorporation of incorrect answers in VARR+ resulted in performance improvements across most datasets. While VARR alone enhanced generation efficiency, VARR+ effectively preserves rationales that help calibrate the probability distribution between correct and incorrect answers, contributing to improved training robustness and stability.
@@ 5.5.1 Identifying Appropriate Units for Removal
Table 1: Analysis across various reasoning reduction units and the application of principled criteria. Each row presents accuracy in the first line, with average generated tokens shown in parentheses in the second line.
In this section, we empirically investigate the necessity of our criteria and demonstrate why sentences are more effective than tokens as reasoning reduction units. While ICoT-SI removes tokens without specific criteria, we first apply VARR+ at the token level (denoted as VARR-Tok) to assess whether tokens can serve as effective reduction units when combined with our criteria. As shown in Table 1, VARR+ applied at the token level achieves an average performance gain of 24.74% compared to ICoT-SI, demonstrating that our principled criteria contribute to robust performance. Furthermore, expanding the reduction units from tokens to sentences (denoted as VARR-Sent) yields an additional performance gain of 15.98% over VARR-Tok. These findings highlight that sentences provide natural and effective boundaries for the reduction process.
@@ 5.5.2 Analyzing the Impact of Sentence Position on Removal Efficacy
Table 2: Performance across 5 different reasoning tasks, evaluated with different sentence position removal. Each row presents accuracy in the first line, with average generated tokens shown in parentheses in the second line.
@@
In Section 3, we demonstrated that gradually removing sentences from random and back positions can degrade model performance. To further explore this finding and assess the robustness of removing sentences from the front position, we conducted experiments with unguided random sentence removal (denoted as No Rule) and applied VARR+ with random position and reverse sentence order (denoted as Random and Back, respectively). As shown in Table 2, unguided random sentence removal resulted in a 25.30% decrease in performance relative to our method, highlighting the critical role of the verbosity evaluation even after selecting sentences as units of reduction. Furthermore, Random and Back strategies exhibited an average 7.50% performance degradation relative to our method. These results further support our observation that earlier sentences in the reasoning path tend to contain more redundancy, and their prioritized removal effectively balances rationale reduction while maintaining reasoning performance.
G8K SQA TQA
Exp-CoT 55.84 74.70 82.94
(122.54) (46.47) (63.46)
RAND-SENT REMOVAL 25.85 72.38 74.14
(63.56) (23.38) (6.84)
FRONT-SENT REMOVAL (1) 38.36 72.96 79.74
(87.43) (24.34) (22.73)
FRONT-SENT REMOVAL (2) 23.50 73.83 76.79
(47.64) (9.97) (8.32)
VARR+ 54.98 78.19 83.45
(100.38) (35.12) (39.17)
Table 3: Performance across different front sentence removal strategies. RAND-SENT-REMOVAL refers to the Random strategy, while ()($\cdot$)( ⋅ ) in the FRONT-SENT REMOVAL indicates the number of enforcedly removed front sentences. Each row presents accuracy in the first line, with average generated tokens shown in parentheses in the second line.
To further validate whether VARR enhances robustness compared to enforced removal of front sentences, we conducted additional experiments where the first and second sentences were enforced to be removed during the initial two training epochs. As shown in Table 3, enforced removal of front sentences leads to substantial performance degradation, while VARR demonstrates its contribution to preserving robustness.
5.5.3 Varying the Warm-up Ratio
@@
Figure 4: Accuracy (barplot) and the average generated token (marker) across various warm-up stages on TriviaQA and CommomsenseQA.
@@
We evaluated our method against various warm-up stages on the TriviaQA and CommonsenseQA datasets. As the duration of the warm-up stages increases, the model becomes more fitted to the non-reduction dataset, which inhibits VARR’s ability to eliminate redundant sentences from the reasoning path. Consequently, as illustrated in Figure 4, longer warm-up periods result in an increase in generated tokens and a decrease in accuracy. These results suggest that 0.1 training steps provide sufficient time to inject reasoning abilities while enabling the systematic reduction of redundant reasoning sentences during the learning process. The zero warm-up configuration demonstrates comparable performance, though 0.1 warm-up steps are slightly better (detailed in Appendix F). Therefore, we set 0.1 as the default setting for the warm-up steps.
@@ 5.5.4 Removal Ratio Analysis
Figure 5: Removal ratio of redundant sentences during training. The y-axis shows the removal ratio, calculated as the number of removed sentences divided by the maximum potential removal sentences (${size(\{r_{i}\})/r(t)}$). The x-axis represents training epochs.
In Figure 5, we analyze the actual amount of rationale sentences removed during training. We examine it by calculating the removal ratio, the proportion of actual removed sentences to the maximum potential removal sentences $r(t)$. Our analysis indicates that not all sentences designated for maximum removal range are always eliminated during the training process. Notably, a significant proportion of redundant sentences are removed in the early stages of training, with fewer sentences being removed as the model progresses through the middle to later stages, thereby stabilizing its training. A similar trend is observed across other datasets, as detailed in Appendix G.
@@ 6 Conclusion
In this work, we propose the sentence-level rationale reduction framework VARR and empirically demonstrate that models trained with non-redundant rationales achieve enhanced efficiency. We address the lack of principled criteria for identifying redundant sentences during training by developing a reduction framework that not only preserves the model’s reasoning capabilities but also reduces the likelihood of generating incorrect answers. Our experiments show that VARR can efficiently handle a diverse range of tasks with fewer generated tokens, without sacrificing its accuracy. This work contributes novel insights to rationale reduction research, contributing to the efficient reasoning elicitation in language models.
@@
While our work provides novel insights into rationale reduction research, our experiments were primarily conducted using a relatively small large language model and limited batch size, constrained by computational costs (i.e., a single A100-80GB GPU). Additionally, for the same reasons, it was not feasible to test the model with datasets featuring long sequences in both queries and rationales (Reddy et al., 2024; Yu et al., 2024a). Nevertheless, given the systematic design principles underlying the VARR/VARR+ frameworks, we believe their effectiveness would extend to larger-scale implementations. Furthermore, we reserve the application of VARR/VARR+ in iterative reasoning path generation and refinement/reflexion-based evaluation discussed in Section 2.1 for future work.
Ethical Considerations
Our work explores how LLMs can maintain their reasoning performance while improving efficiency. To this end, we conducted verbosity-aware rationale reduction (i.e., reasoning sentence pruning)-based CoT fine-tuning, requiring computational resources comparable to standard CoT fine-tuning. Additionally, we used only open-source LLMs and publicly available reasoning datasets with minimal preprocessing using gpt4o-mini’s (OpenAI, 2024) API. Therefore, we do not anticipate significant ethical issues arising from our work. On the contrary, we believe future works could leverage our analysis to reduce computational overhead in CoT inference settings.
Acknowledgements
We sincerely appreciate to Keonwoo Kim and Hyowon Cho for their sharp feedback during writing, Sangyeop Kim, Yukyung Lee, and Minjin Jeon for their continuous encouragement and inspiration during our study sessions, and Eunbi Choi for her invaluable help during the rebuttal process. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2018-0-00584, (SW starlab) Development of Decision Support System Software based on Next-Generation Machine Learning), National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00335873), and National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00217286).